# Using Reinforcement Learning to Train Large Language Models to Explain Human Decisions

### v1

@@ 5 Control Experiments
Using the synthetic choice rate as the reward signal for RL (i.e., replacing $\textbf{p}^{B}$ in Equation 1 with the synthetic choice rate), we find that the RL model quickly learns to generate CoT focused on computing and comparing expected values (see Figure 6b and Appendix D for details). Although risk aversion is no longer necessary to explain the synthetic data, the model still occasionally includes references to it in its CoT. This may be due to residual cues in the prompt suggesting that the task involves human risky choice (see Appendix A.2 for details), even though the actual data had been replaced with synthetic data. Nonetheless, the RL model correctly identifies relevant cognitive mechanisms – such as rational choice theory and all-or-nothing thinking – as important for explaining the synthetic data. These results suggest that RL is capable of adapting its reasoning strategies to match the structure of the training data.
Model control. Additionally, we replicated the main experiment using a smaller and arguably weaker backbone LLM, Gemma-2-2B-Instruct team2024gemma (see Appendix E for details). The same LoRA configuration was applied, resulting in approximately 41.53M trainable parameters, corresponding to 1.56% of the total parameters in the 2B model. The Gemma model was fine-tuned using SFT, Centaur-style SFT, and RL on the same training set as in the main experiment and evaluated on the same held-out test set.
We observe several noteworthy findings. First, RL applied to the Gemma model performs significantly worse than both SFT and Centaur-style SFT in reducing MSE on the test set (see Figure 10 in Appendix E). This is a pattern that contrasts with the learning curves obtained using the stronger Qwen model. Second, analysis of the CoT generated during RL training reveals that the Gemma model fails to compute and compare the expected values of the risky options, a mechanism identified as critical for explaining human behavior in the main experiment (see Appendix E). This absence of key cognitive mechanisms in its CoT likely contributes to the weaker performance of RL on smaller models. Finally, even when RL underperforms, both SFT and Centaur-style SFT are still able to predict human behavior (albeit without explanation) at a level comparable to that of the stronger Qwen model. These suggest that RL and SFT may exhibit different generalization patterns chu2025sft ; wang2025reinforcement .
@@ 6 Discussion
The emergence of reasoning LLMs presents new opportunities for addressing complex problems that traditional, non-reasoning LLMs struggle to solve. This improvement is likely due to the fact that CoT reasoning enhances the expressiveness of Transformer-based models li2024chain . Moreover, RL post-training has been shown to effectively elicit appropriate CoTs from backbone LLMs shao2024deepseekmath ; yue2025does . This process can be seen as analogous to learning to reason like a psychologist: given sufficient capacity, the backbone LLM implicitly encodes a range of psychological theories about human risky choice, and RL serves to surface the most relevant theoretical representations through the generation of CoT reasoning. As a result, we trained an LLM-based psychological theorist capable of verbalizing relevant cognitive mechanisms for explaining and predicting human risky choices.
Failures of RL. During our experiments, we also observed several failures associated with GRPO training. Most notably, applying the original GRPO algorithm shao2024deepseekmath to the choices13k dataset led to mode collapse. By mode collapse, we refer to the phenomenon in which the fine-tuned LLMs converge to generating a single, repetitive reasoning chain that simply computes the expected values of both options and consistently outputs a final prediction of 50% for each option across all choice problems. Note that the choices13k dataset exhibits a modal human choice proportion of 50%, which may have contributed to this degeneracy.
We experimented with several strategies to prevent the model from collapsing to this default prediction. Specifically, we tested alternative reward functions, including one minus the mean squared error, negative cross-entropy loss, and the addition of diversity bonuses incentivizing diverse predictions. However, none of these modifications were effective in mitigating mode collapse in our setting.
The issue was eventually addressed by removing advantage normalization in GRPO, as recommended by liu2025understanding . That is, dividing the centered reward by the standard deviation of group rewards introduces a bias toward question-level difficulty: risky-choice problems with lower standard deviations (i.e., those that are either too easy or too difficult) disproportionately influence policy updates by receiving higher weights. By omitting this normalization step, we were able to recover the main RL results reported above.
The elicitation hypothesis. A prominent hypothesis about RL post-training is that it primarily increases the probability of generating correct outputs by eliciting behaviors already latent in the pretrained LLM lambert2025reinforcement ; yue2025does . In this view, RL post-training does not teach new capabilities but instead amplifies and selects from pre-existing knowledge. Our model control results support this hypothesis: weaker LLMs–particularly those unable to elicit the expected value mechanism for risky choice–struggle to explain human behavior effectively when trained with RL. These results suggest that the success of RL highly depends on the capacity of the backbone LLM.
@@
Limitations and future research. A key limitation of developing cognitive models through CoT lies in the elicitation hypothesis of RL post-training. If this hypothesis holds, RL is unlikely to generate entirely novel theories or cognitive mechanisms. For instance, imagine an LLM pretrained in a world prior to the development of Kahneman and Tversky’s prospect theory kahneman2013prospect . RL alone would likely not discover this theory during post-training. However, in principle, RL can draw on theoretical frameworks beyond psychology to explain psychological phenomena. For example, it may learn to apply concepts from physics or computer science to human behavior, provided such ideas are already embedded in the backbone LLM. Future research could explore how RL facilitates the integration of knowledge across traditional disciplinary boundaries to generate novel explanatory frameworks.
@@
Our work is also constrained by computational resources, as RL is substantially more computationally intensive than SFT. Moreover, our findings suggest that RL and SFT exhibit different generalization patterns. Further research is needed to systematically investigate these differences and understand the conditions under which each method performs the best. An additional open question is when and how to effectively combine RL and SFT to produce more robust and interpretable cognitive models through the CoT of LLMs.
@@
Acknowledgments. JQZ and TLG acknowledge support from the NOMIS Foundation and computing resources provided by Princeton Language and Intelligence. HX acknowledges support from the OpenAI Researcher Access Program. RCW acknowledges funding from a SCIALOG Award (#29079) from the Research Corporation for Scientific Advancement. HX and RCW also acknowledge computing support in part through research cyberinfrastructure resources and services provided by the Partnership for an Advanced Computing Environment at the Georgia Institute of Technology, USA. DA and TLG were supported by ONR MURI N00014-24-1-2748.
@@ A.3 Example prompts for risky options
{mdframed}
Option A offers (1) a 50.0% chance to win \$2.0; (2) a 50.0% chance to win \$0.0.
Option B offers a 100.0% chance to win \$1.0.
Appendix B Cognitive mechanisms identified by Qwen-2.5-7B-Instruct
@@ C.1 Human evaluation of CoT
Participants. We recruited 20 participants via a social media post targeting local universities, inviting volunteers to evaluate completions generated by LLMs. The sample included 12 graduate students, 3 postdoctoral researchers, 3 faculty members, and 2 undergraduate students. Participants represented a range of academic disciplines: 10 from Psychology or Cognitive Science, 7 from Computer Science or Artificial Intelligence, and 3 from Business or Economics. On average, participants reported a self-assessed knowledge of human risky decision-making of 3.5 out of 5. Participants did not receive any monetary payment, and the study lasted up to 30 minutes. All experimental sessions were conducted in May 2025. Informed consent was obtained from all participants (Princeton University IRB number 10859: “Computational Cognitive Science”)
@@
[…]
@@
Due to the large volume of CoTs generated across multiple model checkpoints and risky choice problems, conducting comprehensive human evaluations is constrained by time and labor demands. To address this limitation, we piloted an alternative approach using GPT-4.1 to simulate expert evaluations. Specifically, GPT-4.1 was instructed to role-play as an expert in judgment and decision-making and provide quality ratings of the CoTs. The temperature for GPT-4.1 was fixed at 0. The model was prompted with the following instructions:
{mdframed}
System: You are an expert in judgment and decision-making.
User: As an expert in judgment and decision-making, please evaluate the reasoning and prediction of the following question. Provide a single integer score from 0 to 100 based on the quality of the completion.
@@ C.3 Swapping CoT between backbone and RL models
[…]
@@ E.1 Learning curves
In the model control experiment, we replaced the original Qwen-2.5-7B-Instruct model with the smaller Gemma-2-2B-Instruct model, while keeping the human risky choice data as the target for prediction. We find that SFT and Centaur-style SFT achieve comparable levels of predictive accuracy (see Figure 10). At the final checkpoint, the MSE on the test set is $M=0.0162,SE=0.0006$ for SFT and $M=0.0163,SE=0.0007$ for Centaur-style SFT, with no statistically significant difference between them ($t(2924)=-0.13,p=0.90$). In contrast, RL yields significantly higher MSE at its final checkpoint ($M=0.0526,SE=0.0016$), indicating poorer predictive performance.
Figure 10: Model control experiment. Learning curves on the (a) training and (b) test sets. Backbone LLM is Gemma-2-2B-Instruct. The horizontal axes indicate training epochs, while the vertical axes represent mean squared error (MSE) evaluated on the corresponding dataset. The three post-training strategies compared are supervised fine-tuning (SFT, red), Centaur-style SFT (green), and reinforcement learning using Group Relative Policy Optimization (RL, grey). (c) MSE on the test set at the final checkpoint of each post-training method. Error bars represent ±$\pm 1$ standard error across risky-choice problems.

### v2

@@ 5 Control Experiments
Using the synthetic choice rate as the reward signal for RL (i.e., replacing $\textbf{p}^{B}$ in Equation 1 with the synthetic choice rate), we find that the RL model quickly learns to generate CoT focused on computing and comparing expected values (see Figure 6b and Appendix D for details). Although risk aversion is no longer necessary to explain the synthetic data, the model still occasionally includes references to it in its CoT. This may be due to residual cues in the prompt suggesting that the task involves human risky choice (see Appendix A.2 and D.3 for details), even though the actual data had been replaced with synthetic data. Nonetheless, the RL model correctly identifies relevant cognitive mechanisms (e.g., rational choice theory and all-or-nothing thinking) as important for explaining the synthetic data. These results suggest that RL is capable of adapting its reasoning strategies to match the structure of the training data.
Model control. Additionally, we replicated the main experiment using a smaller and arguably weaker backbone LLM, Gemma-2-2B-Instruct (Team et al., 2024) (see Appendix E for details). The same LoRA configuration was applied, resulting in approximately 41.53M trainable parameters, corresponding to 1.56% of the total parameters in the 2B model. The Gemma model was fine-tuned using SFT, Centaur-style SFT, and RL on the same training set as in the main experiment and evaluated on the same held-out test set.
We observe several noteworthy findings. First, RL applied to the Gemma model performs significantly worse than both SFT and Centaur-style SFT in reducing MSE on the test set (see Figure 12 in Appendix E). This is a pattern that contrasts with the learning curves obtained using the stronger Qwen model. Second, analysis of the CoT generated during RL training reveals that the Gemma model fails to compute and compare the expected values of the risky options, a mechanism identified as critical for explaining human behavior in the main experiment (see Appendix E). This absence of key cognitive mechanisms in its CoT likely contributes to the weaker performance of RL on smaller models. Finally, even when RL underperforms, both SFT and Centaur-style SFT are still able to predict human behavior (albeit without explanation) at a level comparable to that of the stronger Qwen model. These suggest that RL and SFT exhibit different generalization patterns (Chu et al., 2025; Wang et al., 2025).
@@ 6 Discussion
The emergence of reasoning LLMs presents new opportunities for addressing complex problems that traditional, non-reasoning LLMs struggle to solve. This improvement is likely due to the fact that CoT reasoning enhances the expressiveness of Transformer-based models (Li et al., 2024). Moreover, RL post-training has been shown to effectively elicit appropriate CoTs from backbone LLMs (Shao et al., 2024; Yue et al., 2025). This process can be seen as analogous to learning to reason like a psychologist: given sufficient capacity, the backbone LLM implicitly encodes a range of psychological theories about human risky choice, and RL serves to surface the most relevant theoretical representations through the generation of CoT reasoning. As a result, we trained an LLM-based psychological theorist capable of verbalizing relevant cognitive mechanisms for explaining and predicting human risky choices.
RLVR, RLHF, and RLAIF. Our work shows that applying reinforcement learning with verifiable or outcome-based rewards (RLVR) directly to human risky choices can elicit CoTs that explain those decisions at scale. Compared to SFT methods, RL methods have a distinct advantage: the resulting CoTs serve as verbal theories of human choices albeit at the cost of higher computational demands. An alternative to RLVR in the LLM setting is reinforcement learning from human feedback (RLHF; Lambert (2025); Ziegler et al. (2019)). Adapting RLHF to our context would require experts in risky decision-making to compare and evaluate LLM-generated explanations and predictions across different risky choice problems. For instance, scaling up the type of expert judgments illustrated in Figure 7. In practice, this would entail large-scale expert annotations to assess both the model’s CoT explanations and its behavioral predictions, resources that are not currently available.
That said, we acknowledge that a scalable RL alternative may be possible by leveraging stronger LLMs as automated judges (e.g., role-playing expert psychologists), a strategy aligned with reinforcement learning from AI feedback (RLAIF; Bai et al. (2022); Lee et al. (2023)). Of course, scientists could also rely on black-box commercial RL training services (e.g., the RL fine-tuning API offered by OpenAI, n.d.). We chose not to use such services for comparison in the present work due to their lack of transparency, as the details of the underlying RL training are not disclosed.
Training predictive models and then explaining their predictions. As an alternative approach for obtaining interpretable theories of human cognition, researchers can first train predictive models and then apply modern explainability techniques (Agrawal et al., 2020). LLMs can also be used to strengthen the training of such predictive models; for example, synthetic data generated by LLMs has been shown to improve predictive accuracy in economic choice models (Shapira et al., 2024; 2025). This predict-then-explain pipeline is a promising alternative to our RL approach. However, note that the RL method simultaneously optimizes both predictions and explanations using reasoning LLMs. It also avoids the direct use of explainability techniques, which can sometimes be brittle for feature attributions (Bilodeau et al., 2024).
In principle, both approaches aim to uncover explainable and interpretable signals by improving behavioral predictions. The primary difference is that traditional explainability relies on features proposed by human researchers, whereas the RL approach automatically explores features or hypotheses that can explain human data using LLMs. LLM-proposed hypotheses are more automated but depend heavily on the base LLM’s capacity to generate meaningful insights. Human-proposed hypotheses, while slower and less automated, may in some cases be more creative than those generated by LLMs.
@@
Limitations and future research. A key limitation of developing cognitive models through CoT lies in the elicitation hypothesis of RL post-training (Lambert, 2025; Yue et al., 2025). If this hypothesis holds, RL is unlikely to generate entirely novel theories or cognitive mechanisms. For instance, imagine an LLM pretrained in a world prior to the development of Kahneman and Tversky’s prospect theory (Kahneman and Tversky, 2013). RL alone would likely not discover this theory during post-training (see Appendix F for more discussion). However, in principle, RL can draw on theoretical frameworks beyond psychology to explain psychological phenomena. For example, it may learn to apply concepts from physics or computer science to human behavior, provided such ideas are already embedded in the backbone LLM. Future research could explore how RL facilitates the integration of knowledge across traditional disciplinary boundaries to generate novel explanatory frameworks.
@@
Our findings also suggest that RL and SFT exhibit different generalization patterns. Further research is needed to systematically investigate these differences and understand the conditions under which each method performs the best. An additional open question is when and how to effectively combine RL and SFT to produce more robust and interpretable cognitive models through the CoT of LLMs.
@@
Ethics statement
Our human experiment (reported in Appendix C.1) was approved by the Princeton University Institutional Review Board (IRB #10859: ‘Computational Cognitive Science’). All participants provided informed consent prior to the experiment. Participants were asked to evaluate model completions on risky choice problems. We do not anticipate any additional ethical concerns.
Reproducibility statement
We provided complete descriptions of the training methods and associated hyperparameter values in the paper. All model training was conducted using standard Python packages from Hugging Face and vLLM.
Acknowledgments
JQZ and TLG acknowledge support from the NOMIS Foundation and computing resources provided by Princeton Language and Intelligence. HX acknowledges support from the OpenAI Researcher Access Program. RCW acknowledges funding from a SCIALOG Award (#29079) from the Research Corporation for Scientific Advancement. HX and RCW also acknowledge computing support in part through research cyberinfrastructure resources and services provided by the Partnership for an Advanced Computing Environment at the Georgia Institute of Technology, USA. DA and TLG were supported by ONR MURI N00014-24-1-2748.
@@ A.3 Example prompts for risky options
Option A offers (1) a 50.0% chance to win \$2.0; (2) a 50.0% chance to win \$$0.$ Option B offers a 100.0% chance to win \$1.0.
Appendix B Cognitive mechanisms identified by Qwen-2.5-7B-Instruct after RL training
@@ C.1 Human evaluation of CoT
Participants. We recruited 20 participants via a social media post targeting local universities, inviting volunteers to evaluate completions generated by LLMs. The sample included 12 graduate students, 3 postdoctoral researchers, 3 faculty members, and 2 undergraduate students. Participants represented a range of academic disciplines: 10 from Psychology or Cognitive Science, 7 from Computer Science or Artificial Intelligence, and 3 from Business or Economics. On average, participants reported a self-assessed knowledge of human risky decision-making of 3.5 out of 5. Participants did not receive any monetary payment, and the study lasted up to 30 minutes. All experimental sessions were conducted in May 2025. Informed consent was obtained from all participants (Princeton University IRB number 10859: ‘Computational Cognitive Science’)
@@
Results. On average, CoTs generated by the RL model were selected as more reasonable in 61.5% of trials ($SE=5.2\%$) above the chance level of 50% at 0.05 significance level ($t(19)=2.19$, $p<.05$). This suggests that participants found the reasoning produced by the RL model to be more persuasive or coherent than that of the backbone model. While this represents a useful initial step in soliciting expert opinions from researchers in relevant fields, scaling human evaluations to the whole dataset remains a significant challenge.
@@
Due to the large volume of CoTs generated across multiple model checkpoints and risky choice problems, conducting comprehensive human evaluations is constrained by time and labor demands. To address this limitation, we piloted an alternative approach using OpenAI’s GPT-4.1 to simulate expert evaluations. Specifically, GPT-4.1 was instructed to role-play as an expert in judgment and decision-making and provide quality ratings of the CoTs. The temperature for GPT-4.1 was fixed at 0. The model was prompted with the following instructions:
System: You are an expert in judgment and decision-making. User: As an expert in judgment and decision-making, please evaluate the reasoning and prediction of the following question. Provide a single integer score from 0 to 100 based on the quality of the completion.
@@ C.3 Swapping CoT between backbone and RL models
We find that using the RL model’s CoT significantly reduces the backbone model’s prediction error, as measured by MSE (t(2596)=18.11,p<.01t(2596)=-18.11,p<.01). Conversely, using the backbone model’s CoT significantly increases the RL model’s prediction error ($t(2597)=20.65,p<.01$). These results suggest that the RL model generates higher-quality CoTs that contribute more effectively to accurate final predictions than those produced by the backbone model.
@@
C.4 Testing the CoTs Generated by SFT and Centaur Models
When an LLM is SFT-trained on risky choices, it may also learn the behavioral policy reflected in the risky-choice data (Betley et al., 2025). For example, if the choice data are risk-seeking, the finetuned LLM will exhibit more risk-seeking behavior as well. Therefore, to further test whether the CoTs elicited using RL are more informative than those produced via SFT, we also elicited CoTs from SFT models and evaluate whether these SFT-generated CoTs are effective at predicting human risky choices. Specifically, we prompted both the SFT and Centaur models (i.e., the final SFT and Centaur checkpoints) with the RL prompt that encourages step-by-step reasoning before providing a behavioral prediction in JSON format. We then removed the final JSON prediction from the model completions and fed the remaining CoT, without the JSON output, to the backbone LLM – a procedure similar to the CoT-swapping experiment reported in Table 2.
We find that neither SFT-generated nor Centaur-generated CoTs improve the backbone model’s predictions of human risky choices. The MSE of the backbone LLM’s predictions using SFT-generated CoTs is M=0.0785(SE=0.0023)M=0.0785~(SE=0.0023), and the MSE using Centaur-generated CoTs is M=0.0728(SE=0.0019)M=0.0728~(SE=0.0019). Both values are comparable to the performance of the backbone LLM using its original CoT (see the top-left cell of Table 2). These results suggest that although SFT and Centaur models can achieve comparable behavioral predictions without CoTs, the CoTs generated by these models do not help improve the predictions of the original LLM. This highlights the particular need for RL post-training, beyond SFT and Centaur-style SFT, to produce more effective CoTs.
Finally, we examined the evolution of SFT-generated and Centaur-generated CoTs over the course of training. Similar to the analysis in Figure 6, where we used GPT-4.1 to summarize CoTs generated by the RL models, we also prompted GPT-4.1 with the SFT-generated and Centaur-generated CoTs (see Figure 9). Compared to the CoTs elicited through RL training in Figure 6, both SFT and Centaur-style SFT (neither of which are explicitly trained to produce CoTs) showed no noticeable changes in CoT frequencies across training epochs. We do observe an upward trend in the usage of “expected value” in the SFT-generated CoTs (Figure 9a), whereas its usage decreases in the Centaur-generated CoTs (Figure 9b). This suggests that even though SFT and Centaur-style SFT yield similar behavioral predictions, the CoTs they generate can differ.
Figure 9: Evolution of Chain-of-Thoughts (CoTs) generated by the SFT and Centaur-style SFT models over training epochs. (a) SFT-generated CoTs. (b) Centaur-generated CoTs.
@@ D.1 Learning curves
D.3 Additional Data Control Experiments
To further verify that the CoTs elicited from LLMs through RL training adapt to the structure of the training dataset, we conducted two additional control experiments using (i) a Random Dataset, consisting of randomly generated choice rates (i.e., samples from $U[0,1]$), and (ii) a Complexity Aversion Dataset, in which choice rates reflect only complexity aversion. Specifically for (ii), we define complexity based on the number of outcomes in each gamble. For example, if Option A had one outcome and Option B had two, the synthetic data specified that $2/3$ of people would prefer Option A (i.e., $\frac{\text{num\_outcomes\_B}}{\text{num\_outcomes\_A}+\text{num\_outcomes\_B}}$), and $1/3$ would choose Option B (i.e., $\frac{\text{num\_outcomes\_A}}{\text{num\_outcomes\_A}+\text{num\_outcomes\_B}}$) regardless of the underlying probabilities or payoff values. In essence, the only decision rule was to probabilistically prefer the option with fewer outcomes.
We then trained the Qwen-2.5-7B-Instruct model using the same RL setup in the main text. When trained on the Random Dataset, the model collapsed to consistently predicting 50% choice rates, and its CoT generations became nonsensical, often degenerating into multilingual gibberish. In contrast, when trained on the Complexity Aversion Dataset, the learning curve showed a steady increase in correctness reward toward 0.97 (with 1 as the maximum), indicating that the model has successfully learned to predict complexity-averse behavior. Moreover, the resulting CoTs generated by the RL model more frequently referenced concepts such as “simplicity” and “complexity” (e.g., “Complexity: Option B is significantly more complex and involves a higher number of possible outcomes. People often avoid options with too many uncertainties and choices, which can lead to decision fatigue and heuristic shortcuts.”).
However, as shown in Figure 11a, the CoTs did not exclusively focus on complexity, likely due to residual influence from the original prompt, which framed the task as explaining human risky choice. This phenomenon mirrors our earlier data control experiment in which synthetic data were generated by an expected-value maximization model. However, the relative rank of thoughts that mention complexity aversion steadily increases over the course of RL training, rising from around 16th place to around 6th place (see Figure 11b). Taken together, these additional control experiments further support the flexibility and robustness of our RLVR-based approach in eliciting sensible CoT explanations from LLMs that adapt to the structure of the training data.
Figure 11: Proportions of cognitive mechanisms identified in the RL models’ thoughts across training epochs in the additional data control experiment, where the synthetic dataset reflects complexity-aversion behavior. (a) The eight most frequently used mechanisms are shown. (b) The relative rank of complexity-aversion usage in the CoTs increases over the course of RL training.
@@ E.1 Learning curves
In the model control experiment, we replaced the original Qwen-2.5-7B-Instruct model with the smaller Gemma-2-2B-Instruct model, while keeping the human risky choice data as the target for prediction. We find that SFT and Centaur-style SFT achieve comparable levels of predictive accuracy (see Figure 12). At the final checkpoint, the MSE on the test set is $M=0.0162,SE=0.0006$ for SFT and $M=0.0163,SE=0.0007$ for Centaur-style SFT, with no statistically significant difference between them (t(2924)=0.13,p=0.90t(2924)=-0.13,p=0.90). In contrast, RL yields significantly higher MSE at its final checkpoint ($M=0.0526,SE=0.0016$), indicating poorer predictive performance.
Figure 12: Model control experiment. Learning curves on the (a) training and (b) test sets. Backbone LLM is Gemma-2-2B-Instruct. The horizontal axes indicate training epochs, while the vertical axes represent mean squared error (MSE) evaluated on the corresponding dataset. The three post-training strategies compared are supervised fine-tuning (SFT, red), Centaur-style SFT (green), and reinforcement learning using Group Relative Policy Optimization (RL, grey). (c) MSE on the test set at the final checkpoint of each post-training method. Error bars represent $\pm 1$ standard error across risky-choice problems.